# Plurals: A System for Guiding LLMs Via Simulated Social Ensembles

### v5

@@ 5. Case Studies
Figure 4. In three experiments, both zero-shot and Plurals simulated focus groups tried to create output compelling to specific audiences. Plurals simulated focus group output was chosen by an online sample of the relevant audiences over zero-shot. See Appendix Table 3 for multilevel regressions.
@@
We provide several preliminary empirical results (Table 1). Case Studies 1 and 2 are meant as mechanistic fidelity checks. We show that the system does what we are claiming it does. Case Studies 3-5 are efficacy tests. We show that our system outperforms a standard zero-shot (and zero-shot chain-of-thought) LLM approach. Case Study 6 is a preliminary analysis of how this system can be used for ethical guardrails. All human subject experiments received prior IRB approval from our university and met power requirements [^6] Two-tailed exact binomial test parameters (observed proportion vs. 0.5): $g=0.1,\beta=0.8,\alpha=0.05$, computed using G*Power 3.1.; Note that exact binomial tests do not rely on asymptotic assumptions..
@@
In Case Study 1, we show that using intersectional ANES personas (i.e: combining ideology with demographic variables) results in more response diversity than prompting with only-ideology personas (“You are a liberal”), suggesting this multi-attribute persona method can reduce homogenization. In Case Study 2, we show that Agents are correctly applying a subset of combination instructions from our deliberation templates—suggesting that combination instructions can steer LLM deliberation.
@@
In Case Studies 3-5, we used zero-shot and Plurals simulated social ensembles to create output aimed at resonating with specific audiences. Plurals output was chosen as more compelling by the relevant Prolific audience for both conservatives (Study 3) and liberals (Study 4, Study 5). See Appendix Table 3 for multilevel models and a pooled analysis across studies. In Case Study 6, we discuss how Plurals can facilitate custom ethical guardrails with a preliminary case study. We discuss the ethical implications of studies 3-6 in more detail in section 7.
@@
We discussed how intersectional personas from government datasets should lead to less homogenizing output than single-attribute personas. Responses for a set of prompts corresponding to different liberals (“You are a liberal and $X=x$ and $Y=y$…”) should logically have more diversity than applying the same single-ideology prompt (“You are a liberal.”). Here we show this empirically. Our ANES persona method for political ideologies generates more diverse responses than prompting an LLM with only ideology instructions in 100% of Claude Sonnet comparisons and 95% of GPT-4o comparisons. This is almost true by definition, so methodology and analysis are in Appendix C.
@@
We first collected 2024 ballot initiatives from the website Ballotpedia. We then randomly sampled 30 of the 137 ballot measures for which we could scrape both a short description and a more detailed explanation to turn into a prompt (Appendix D). We then generated two-cycle debates for each ballot initiative under rational and emotional conditions, differing only in one line of combination instructions [^7] Rational: “Give more weight to rational arguments rather than emotional ones.”; Emotional: “Give value to emotional forms of communication, such as narrative, rhetoric, testimony, and storytelling.”. We used the final response from each debate for annotation, with agents randomly assigned to be GPT-4o, GPT-4 Turbo, or Claude Sonnet. See Appendix D for full combination instructions.
@@
We recruited 20 participants from Prolific who completed more 100 tasks and had a 98%+ approval rating. Participants were paid \$2, based on an anticipated study duration of 7 minutes (\$17/hr). After providing informed consent, each participant viewed 10 pairs of responses (rational, emotional) for different ballot measures. We randomly assigned participants to identify either the rational or emotional condition across their 10 trials. We randomized both the order of condition presentation within each pair and the sequence of ballot measures. See Appendix D for task wording.
@@ Results
5.3. Efficacy: Conservatives preferred solar panel ideas from a simulated focus group of conservatives over zero-shot.
@@
Combining ANES personas, ensembles, and Moderators, we tested whether a “simulated focus group” yields ideas resonant with the relevant Prolific audience. Specifically, the aim of this study was to generate descriptions of solar panel companies that conservatives would buy from. We generated descriptions under two conditions—zero-shot, or a simulated focus group of ANES conservatives. In the latter, we queried 10 simulated ANES conservatives on what they would want in a solar panel company, and then a Moderator proposed an idea based on this simulated feedback. We then had conservatives on Prolific evaluate solar panel company ideas and found those generated from the simulated focus group were preferred over zero-shot ideas in 88% of cases. This experiment used GPT-4o. See Appendix F for code.
@@
In earlier pilots, we found that simply prompting LLMs to generate ideas for a solar panel company for conservatives resulted in outputs that were highly ideological (e.g., emphasizing being founded by a veteran). This was despite instructions like “be very specific” that we maintained for this study. However, when LLMs simulated specific conservatives who were asked what product details they would want in a solar panel company, few of the product details were ideological. Hence, our intuition was that this focus group would surface concerns relevant to actual conservatives (e.g: rural weather) as a function of the non-ideological aspects of the conservative ANES personas. More generally, personalization (incorporating details about a user into messaging) increases the persuasiveness of LLM generations (simchon_persuasive_2024, ). Querying simulated personas can be thought of as a synthetic kind of “personalization”.
Human Experiment
We recruited 20 conservative participants from Prolific using Prolific’s screening tool [^8] Participants were asked: “Where would you place yourself along the political spectrum?” and allowable options were: Conservative, Moderate, Liberal, other, N/A. We applied additional filters to ensure participants lived in the United States, were above 18, and had a 98% approval rating. Participants were paid \$1.50 for an expected duration of 6 minutes (\$15/hr). After providing informed consent, participants answered a commitment check (qualtrics_using_2022, ) affirming they would provide high-quality data. Then, for 15 trials, participants were shown pairs of solar panel company ideas generated under both zero-shot and the simulated focus group. Participants were asked, “Supposing that you were going to make a purchase from a solar panel company, which company would you choose?” We randomized the presentation order of condition responses and sequence of idea pairs.
Measures
We conducted exact two-tailed binomial tests on whether the proportion of times the simulated focus group option was chosen differed from chance.
Results
We find that the focus group output was chosen over the zero-shot output in 88% of cases (95% CI = [84%, 91%]), binomial $p<0.001$, Figure 4.
5.4. Efficacy: Liberals preferred charter school ideas from a simulated focus group of liberals over zero-shot.
@@
We conducted a follow-up experiment to the solar panel experiment. Here, the goal was to generate descriptions of charter schools that liberal parents would send a child to. Using a similar setup—and evaluations from liberals with children—we found those descriptions generated from the simulated focus group were preferred over zero-shot chain of thought (CoT) ideas in 69% of cases. This experiment used Claude Sonnet. See Appendix G for materials and code.
@@
In the zero-shot condition, we generated a charter school idea using a CoT prompt. In the Plurals (DAG) condition, we also started with a CoT idea. But then this initial idea was fed to three simulated liberal parents, who offered separate critiques of the idea. Then a default Agent executed a variant of the initial CoT prompt, taking into account critiques of the initial idea. We generated 15 pairs of zero-shot ideas and DAG ideas. See Appendix G for code. This experiment differed from the previous experiment in two ways. We used a CoT prompt for the zero-shot generation since this may be a more difficult baseline. We also employed a “critique and revise” setup similar to the idea behind constitutional AI (CAI) (bai_constitutional_2022, ).
Human Experiment
We recruited 20 liberal parents from Prolific, using Prolific’s screening tool [^9] Participants were asked: “Where would you place yourself along the political spectrum?” and allowable options were: Conservative, Moderate, Liberal, other, N/A. Participants were also asked: “Do you have any children?” and allowable options were Yes, No.. We applied additional filters to ensure participants lived in the United States, were above 18, and had a 98% approval rating. Participants earned \$1.75 for an expected duration of 7 minutes (\$15/hr). After providing informed consent, participants answered a commitment check (qualtrics_using_2022, ). We then presented a brief passage on charter schools adapted from Wikipedia (noauthor_charter_2024, ), followed by a comprehension check (Appendix G) of this passage. For 15 trials, participants chose between pairs of charter school ideas generated under zero-shot and simulated focus group conditions, answering, “Supposing you were sending a child to a charter school, which would you choose?” We randomized the presentation order of condition responses and sequence of idea pairs. See Appendix G for the comprehension check.
Measures
We conducted exact two-tailed binomial tests on whether the proportion of times the simulated focus group option was chosen differed from chance.
Results
We find that the focus group output was chosen over the zero-shot output in 69% of cases, (95% CI = [63%, 74%]), binomial $p<0.001$, Figure 4.
5.5. Efficacy: Liberals preferred homeless shelter ideas from a simulated focus group of liberals over zero-shot.
@@
We conducted a third efficacy experiment that was motivated by “NIMBYism” (Not in My Backyard)—the phenomena of citizens supporting policies in the abstract but not in their specific neighborhoods (davidson_beyond_2014, ; young_necessary_2012, ; lyon-callo_making_2001, ). Here, the goal was to generate proposals for homeless shelters—which are a frequent target of NIMBYism (young_necessary_2012, ; lyon-callo_making_2001, )—that liberals would find compelling. Using a similar setup to the previous experiment, we find that ideas generated from the simulated focus group were preferred over zero-shot ideas in 66% of trials. This experiment used Claude Sonnet. See Appendix H for code.
@@
Human Experiment
We recruited 20 liberals from Prolific who lived in the United States, were above 18, and had a 98% approval rating. Participants were paid \$1.75 for an expected duration of 7 minutes (\$15/hr). After providing informed consent, participants answered a commitment check (qualtrics_using_2022, ) and then engaged in 10 trials. In each trial, participants were shown pairs of homeless shelter proposals generated under both zero-shot and the simulated focus group and were asked, “Consider two proposals for a homeless shelter in your neighborhood. Which of these proposals would be more compelling to you?” We randomized the presentation order of condition responses and sequence of idea pairs.
Measures
We conducted exact two-tailed binomial tests on whether the proportion of times the simulated focus group option was chosen differed from chance.
Results
Plurals output was chosen over zero-shot in 66% of cases, (95% CI = [60%, 73%]), binomial $p<0.001$, Figure 4.
@@ 5.6. Moderation: Using Plurals for LLM Guardrails
Case studies 3-5 demonstrate Plurals’ ability to create output that resonates with audiences more than zero-shot approaches. However, depending on the use, this capability raises ethical concerns—which we discuss more extensively in section 7. Here, we present a case study on steerable Moderators as an initial exploration of how Plurals abstractions can create ethical guardrails. We show that Moderators can be steered to accept or reject requests, based on specific values they are initialized with, at 91% accuracy.
@@
While previous experiments showed how Moderators can improve participants’ outputs, Moderators can also decide whether to proceed with synthesis or reject requests outright. Consider a structure, for instance, where Agents deliberate and a Moderator decides whether to pass on this output to users. Or consider a system where the subject of multi-agent deliberation is whether to process the request, in the first place. These are all examples of “steerable moderation”. This case study provides initial insights into how one could use Plurals for steerable moderation, laying the groundwork for future research on Plurals deliberation for guiding LLM abstentions (an area we plan to explore in future work).
@@
We began with Abercrombie et al.’s (abercrombie_collaborative_2024, ) typology of AI, algorithmic, and automation harms. We selected two specific harms—environmental and physical harms. For each harm, we crafted three user prompts that would trigger concerns in one category but not the other (Appendix I), testing the Moderator’s ability to discriminate between tasks based on their specific value sets. We initialized Moderators with specific value sets using a CoT system prompt that incorporated Abercrombie et al.’s language around typology definitions (Appendix I), instructing Moderators to abstain from processing tasks if and only if the task conflicted with their assigned values. Using GPT-4o, we conducted 30 iterations per (task, value) combination, resulting in 360 total annotations. In each iteration, a Moderator decides whether to accept or reject the given task.
@@
Our primary metric was abstention accuracy. We defined a decision as “accurate” when a moderator abstains if and only if the task violates its assigned value. We used two-tailed binomial tests to determine if the accuracy differed significantly from chance.
@@ 6. Limitations and Future Work
Because large language models are trained on specific datasets and in specific ways, there are logical limits to the extent to which they can be steered. They may, for example, internalize distinct priors (ashkinaze_seeing_2024, ). In some cases, prompting can help mitigate this fixedness. Anecdotally, through development, we found that models adhered more to ANES personas when an instruction included language such as avoiding being “overly polite”. (Relatedly, research finds LLMs tend to be sycophantic (ranaldi_when_2024, ; sharma_towards_2023, ), likely a result of preference alignment (sharma_towards_2023, ).) However, it is not obvious beforehand the extent to which LLMs can be steered to complete tasks. A lack of steerability may limit the model’s ability to simulate different perspectives.
@@
Separate from steerability is the question of how faithful LLM personas are. Prior research suggests LLMs can effectively model personas (argyle_out_2023, ; gao_peacok_2023, ; benharrak_writer-defined_2024, ; milicka_large_2024, ; li_steerability_2024, ) while other research shows LLM personas fail to replicate desired behaviors (von_der_heyde_assessing_2023, ; kovac_stick_2024, ; dong_i_2024, ). Our ANES implementation is based on (argyle_out_2023, ), where Argyle et al. (argyle_out_2023, ) showed this method produces accurate responses when measured against participant responses from ANES. Of course, there are more ways to generate personas than via government datasets. For example, Li et al. (li_steerability_2024, ) find constructing personas from specific custom behaviors is effective. In future iterations, we plan on adding additional persona-generation methods. We also note that our package can be used in the absence of personas. For example, users may be interested in customizing information-sharing Structures and using models without personas.
@@
We face two distinct challenges regarding LLM personas: an empirical question about their fidelity and a larger methodological question about the necessary level of fidelity for utility. For instance, human evaluations of semantic embeddings do not correlate with downstream task performance (chiu_intrinsic_2016, ; bakarov_survey_2018, ). Similarly, we propose that researchers consider the purpose of personas. If the end goal is replacements for people, even setting aside the significant ethical concerns, they would require very high fidelity. But if personas are used as tools to augment human decision-making in specific contexts, the required fidelity (and even how to measure fidelity) likely varies by task.
@@
There is still (relatively) little research on how best to steer large language models to incorporate new information from prior Agents optimally (zhang_chain_2024, ). For example, it is possible that a prior Agent’s response degrades the performance of a future Agent. These questions are highly relevant as practitioners are increasingly using retrieval-augmented generation (RAG) along with LLMs (gao_retrieval-augmented_2023, ). Our package can serve as a useful testbed for researchers who are studying how best to combine and filter new information to complete tasks. In human diffusion, initial behavior has a large effect on cascades (salganik_leading_2008, ; muchnik_social_2013, ). Plurals can be used to understand: What structures and combination instructions minimize undesirable Agent-based cascades (ju_flooding_2024, )?
@@
We chose ANES as an initial persona-generation dataset due to its use in prior work (argyle_out_2023, ), inclusion of political variables, and updating frequency. Nonetheless, ANES is just one possible generator of data-driven personas and is limited to the United States, does not represent non-citizens, and is heavily focused on demographic and political variables. In future iterations, we plan on adding orthogonal datasets.
@@
Our efficacy studies showed our system is an improvement over zero-shot but this does not necessarily mean it is helpful in general—just that it beats a baseline. We also did not systematically explore the efficacy of Plurals. Future work can explore different baselines such as expert-crafted messages, different structures, and different simulated Agents. Mechanistic fidelity case studies correspond to personas and (a slice of) combination instructions. Future work can explore the fidelity of more components. Our steerable moderation case study is a simplified proof of concept since many tasks do not cleanly violate just one principle and not others. Future work can more thoroughly evaluate whether Plurals can be steered to accurately abstain, based on user-defined values. These case studies are a preliminary exploration of Plurals.

### v6

@@ 5. Case Studies
Figure 4. In three experiments, both zero-shot and Plurals simulated focus groups tried to create output compelling to specific audiences. Plurals simulated focus group output was chosen by an online sample of the relevant audiences over zero-shot. See SM Table 1 for multilevel regressions.
[…]
This figure summarizes the results of three experiments. In three experiments, both zero-shot and Plurals simulated focus groups tried to create output compelling to specific audiences. A bar chart shows Plurals simulated focus group output was chosen by an online sample of the relevant audiences over zero-shot.
@@
We provide several preliminary empirical results (Table 1). Case Studies 1 and 2 are mechanistic fidelity checks. We show that the system does what we are claiming it does. Case Studies 3-5 are efficacy tests. We show that our system outperforms a standard zero-shot (and zero-shot chain-of-thought) LLM approach. Case Study 6 is a preliminary analysis of how this system can be used for ethical guardrails. All human subject experiments received prior IRB approval from our university and met power requirements [^6] Two-tailed exact binomial test parameters (observed proportion vs. 0.5): $g=0.1,\beta=0.8,\alpha=0.05$, computed using G*Power 3.1.; Note that exact binomial tests do not rely on asymptotic assumptions..
Rationale & Implications for Mechanistic Fidelity Experiments.
@@
In Case Study 1, we show that using intersectional ANES personas (i.e.: combining ideology with demographic variables) results in more response diversity than prompting with only-ideology personas (“You are a liberal”), suggesting this multi-attribute persona method can reduce homogenization. In Case Study 2, we show that Agents can apply a subset of our first- and second-generation deliberation ideals correctly. We chose these specific combination instructions—instructing agents to emphasize either rational or emotional arguments—because they likely have broad applicability. Case Study 2 provides proof-of-concept that combination instructions can correctly steer LLM deliberations.
Rationale & Implications for Efficacy Experiments.
@@
In Case Studies 3-5, we used zero-shot and Plurals simulated social ensembles to create output aimed at resonating with specific audiences. Plurals output was chosen as more compelling by an online sample of the relevant audience for both conservatives (Study 3) and liberals (Study 4, Study 5). These case studies show that relative to non-Plurals LLM generation, Plurals is more effective at resonating with target audiences. We discuss the ethical implications of system efficacy in section 7.
To evaluate efficacy, we (1) evaluated our system on both conservatives and liberals and (2) chose polarized domains where individual preferences may be more nuanced than political ideology, alone. Solar panel adoption is illustrative: Republicans are less supportive of solar panels in the abstract (Kennedy, 2024) but are highly responsive to material incentives in practice (Dokshin and Gherghina, 2024). Liberals are less supportive of charter schools (B. Henderson et al., 2019) but parents’ educational priorities may not be purely ideological. Even for communities that would in theory be ideologically accepting, homeless shelters are frequently the target of “Not in My Backyard” (NIMBY-ism) (Orr et al., 2024). For two experiments, we strengthened our baseline by using chain-of-thought generation. We chose vanilla zero-shot and chain-of-thought as baselines since fine-tuned and few-shot models might require examples a developer does not have. Plurals significantly outperformed baselines in all experiments. See section 6 for limitations and future directions.
Rationale & Implications for Moderation Experiment.
In Case Study 6, we discuss how Plurals can facilitate custom ethical guardrails with a preliminary case study. This case study shows Plurals may be able to reject requests based on custom values, an area we plan to build on in future work.
@@
We discussed how intersectional personas from government datasets should lead to less homogenizing output than single-attribute personas. Responses for a set of prompts corresponding to different liberals (“You are a liberal and $X=x$ and $Y=y$…”) should logically have more diversity than applying the same single-ideology prompt (“You are a liberal.”). Here we show this empirically. Our ANES persona method for political ideologies generates more diverse responses than prompting an LLM with only ideology instructions in 100% of Claude Sonnet comparisons and 95% of GPT-4o comparisons. This is almost true by definition, so methodology and analysis are in Appendix B.
@@
We first collected 2024 ballot initiatives from the website Ballotpedia. We then randomly sampled 30 of the 137 ballot measures for which we could scrape both a short description and a more detailed explanation to turn into a prompt (Appendix C). We then generated two-cycle debates for each ballot initiative under rational and emotional conditions, differing only in one line of combination instructions [^7] Rational: “Give more weight to rational arguments rather than emotional ones.”; Emotional: “Give value to emotional forms of communication, such as narrative, rhetoric, testimony, and storytelling.”. We used the final response from each debate for annotation, with agents randomly assigned to be GPT-4o, GPT-4 Turbo, or Claude Sonnet. See Appendix C for full combination instructions.
@@
We recruited 20 participants from Prolific who completed more 100 tasks and had a 98%+ approval rating. Participants were paid \$2, based on an anticipated study duration of 7 minutes (\$17/hr). After providing informed consent, each participant viewed 10 pairs of responses (rational, emotional) for different ballot measures. We randomly assigned participants to identify either the rational or emotional condition across their 10 trials. We randomized both the order of condition presentation within each pair and the sequence of ballot measures. See Appendix C for task wording.
@@ Results
5.3. Efficacy: Simulated focus groups create compelling output.
Common Experiment Setup.
We conducted three experiments to test whether Plurals’ simulated focus groups could create output that resonates with specific audiences. All three efficacy experiments followed a similar procedure. We first generated output via zero-shot and a Plurals simulation of an audience. We then recruited members of each audience through Prolific (additional filters: 98%+ approval rating, lived in the United States, were above 18). Compensation was set to over \$15/hr for each experiment. After providing informed consent, participants completed a commitment check (Qualtrics, 2022). Then, participants viewed pairs of responses (zero-shot vs. Plurals) in a masked and randomized order and selected which they found more compelling. We conducted two-tailed binomial tests on whether Plurals was chosen at a rate that differed from chance.
5.4. Efficacy Experiment 1: Conservatives preferred solar panel ideas from a simulated focus group of conservatives over zero-shot.
@@
Using GPT-4o, we generated solar panel company descriptions that would appeal to conservatives. A simulated focus group of conservatives generated ideas that Prolific conservatives preferred over zero-shot ideas in 88% of cases. See SM section 3 for materials.
@@
In earlier pilots, we found that simply prompting LLMs to generate ideas for a solar panel company for conservatives resulted in outputs that were highly ideological (e.g., emphasizing being founded by a veteran). This was despite instructions like “be very specific” that we maintained for this study. However, when LLMs simulated specific conservatives who were asked what product details they would want in a solar panel company, few of the product details were ideological. Hence, our intuition was that this focus group would surface concerns relevant to actual conservatives (e.g.: rural weather) as a function of the non-ideological aspects of the conservative ANES personas. More generally, personalization (incorporating details about a user into messaging) increases the persuasiveness of LLM generations (Simchon et al., 2024). Querying simulated personas can be thought of as a synthetic kind of “personalization”.
Human Evaluation
We recruited 20 conservative participants from Prolific using Prolific’s screening tool [^8] Participants were asked: “Where would you place yourself along the political spectrum?” and allowable options were: Conservative, Moderate, Liberal, other, N/A who engaged in 15 trials each. In each trial, participants were shown pairs of solar panel company ideas generated under both zero-shot and the simulated focus group. Participants were asked, “Supposing that you were going to make a purchase from a solar panel company, which company would you choose?” Plurals output was chosen in 88% of cases (95% CI = [84%, 91%]), binomial $p<0.001$, Figure 4.
5.5. Efficacy Experiment 2: Liberals preferred charter school ideas from a simulated focus group of liberals over zero-shot.
@@
Using Claude Sonnet, we conducted a follow-up experiment to the solar panel experiment. Here, the goal was to generate descriptions of charter schools that liberal parents would send a child to. The Plurals approach outperformed zero-shot chain-of-thought (CoT) generation, with liberals preferring Plurals output in 69% of cases. See SM section 4 for materials.
@@
In the zero-shot condition, we generated a charter school idea using a CoT prompt. In the Plurals (DAG) condition, we also started with a CoT idea. But then this initial idea was fed to three simulated liberal parents, who offered separate critiques of the idea. Then a default Agent executed a variant of the initial CoT prompt, taking into account critiques of the initial idea. We generated 15 pairs of zero-shot ideas and DAG ideas. This experiment differed from the previous experiment in two ways. We used a CoT prompt for the zero-shot generation as a more difficult baseline. We also employed a “critique and revise” setup similar to the idea behind constitutional AI (CAI) (Bai et al., 2022).
Human Evaluation
We recruited 20 liberal parents from Prolific, using Prolific’s screening tool [^9] Participants were asked: “Where would you place yourself along the political spectrum?” and allowable options were: Conservative, Moderate, Liberal, other, N/A. Participants were also asked: “Do you have any children?” and allowable options were Yes, No. who engaged in 15 trials each. Participants first read a brief passage on charter schools adapted from Wikipedia (Wikipedia, 2024), followed by a comprehension check. For each trial, participants chose between pairs of charter school ideas generated under zero-shot and simulated focus group conditions, answering, “Supposing you were sending a child to a charter school, which would you choose?” Plurals output was chosen in 69% of cases, (95% CI = [63%, 74%]), binomial $p<0.001$, Figure 4.
5.6. Efficacy Experiment 3: Liberals preferred homeless shelter ideas from a simulated focus group of liberals over zero-shot.
@@
We conducted a third efficacy experiment that was motivated by “NIMBYism” (Not in My Backyard)—the phenomena of citizens supporting policies in the abstract but not in their specific neighborhoods (Scally and Tighe, 2015; Dear, 1992). Here, the goal was to generate proposals for homeless shelters—which are a frequent target of NIMBYism (Orr et al., 2024)—that liberals would find compelling. Using Claude Sonnet, our simulated focus group generated proposals that liberals preferred over zero-shot ideas in 66% of trials. See SM section 5 for materials.
@@
Human Evaluation
We recruited 20 liberals from Prolific who engaged in 10 trials each. For each trial, participants were shown pairs of homeless shelter proposals generated under both zero-shot and the simulated focus group and were asked, “Consider two proposals for a homeless shelter in your neighborhood. Which of these proposals would be more compelling to you?”. Plurals output was chosen in 66% of cases, (95% CI = [60%, 73%]), binomial $p<0.001$, Figure 4.
@@ 5.7. Moderation: Using Plurals for LLM Guardrails
Case Studies 3-5 demonstrate Plurals’ ability to create output that resonates with audiences more than zero-shot approaches. However, depending on the use, this capability raises ethical concerns—which we discuss more extensively in section 7. Here, we present a case study on steerable Moderators as an initial exploration of how Plurals abstractions can create ethical guardrails. Moderators can be steered to accept or reject requests, based on specific values they are initialized with, at 91% accuracy.
@@
While previous experiments showed how Moderators can improve participants’ outputs, Moderators can also decide whether to proceed with synthesis or reject requests outright. Consider a structure, for instance, where Agents deliberate and a Moderator decides whether to pass on this output to users. Or consider a system where the subject of multi-agent deliberation is whether to process the request. These are examples of “steerable moderation”. This case study provides initial insights into how one could use Plurals for steerable moderation, laying the groundwork for future research on Plurals deliberation for guiding LLM abstentions (an area we plan to explore in future work).
@@
We began with Abercrombie et al.’s (Abercrombie et al., 2024) typology of AI, algorithmic, and automation harms. We selected two specific harms—environmental and physical harms. For each harm, we crafted three user prompts that would trigger concerns in one category but not the other (Appendix D), testing the Moderator’s ability to discriminate between tasks based on their specific value sets. We initialized Moderators with specific value sets using a CoT system prompt that incorporated Abercrombie et al.’s language around typology definitions (Appendix D), instructing Moderators to abstain from processing tasks if and only if the task conflicted with their assigned values. Using GPT-4o, we conducted 30 iterations per (task, value) combination, resulting in 360 total annotations. In each iteration, a Moderator decides whether to accept or reject the given task.
@@
Our primary metric was abstention accuracy, defined as abstaining if and only if the task violates the Moderator’s assigned value. We used two-tailed binomial tests to determine if the accuracy differed from chance.
@@ 6. Limitations and Future Work
Because large language models are trained on specific datasets and in specific ways, there are logical limits to the extent to which they can be steered. They may, for example, internalize distinct priors (Ashkinaze et al., 2024a). In some cases, prompting can help mitigate this fixedness. Anecdotally, through development, we found that models adhered more to ANES personas when an instruction included language such as avoiding being “overly polite”. (Relatedly, research finds LLMs tend to be sycophantic (Ranaldi and Pucci, 2024; Sharma et al., 2023), likely a result of preference alignment (Sharma et al., 2023).) However, it is not obvious beforehand the extent to which LLMs can be steered to complete tasks. A lack of steerability may limit the model’s ability to simulate different perspectives.
@@
Separate from steerability is the question of how faithful LLM personas are. Prior research suggests LLMs can effectively model personas (Argyle et al., 2023b; Gao et al., 2023a; Milička et al., 2024; Li et al., 2024) while other research shows LLM personas fail to replicate desired behaviors (von der Heyde et al., 2023; Kovač et al., 2024; Dong et al., 2024). Our ANES implementation is based on (Argyle et al., 2023b), where Argyle et al. showed this method produces accurate responses when measured against participant responses from ANES. Of course, there are more ways to generate personas than via government datasets. In future iterations, we plan on adding additional persona-generation methods. We also note that our package can be used in the absence of personas. For example, users may be interested in customizing information-sharing Structures and using models without personas.
@@
We face two distinct challenges regarding LLM personas: an empirical question about their fidelity and a larger methodological question about the necessary level of fidelity for utility. For instance, human evaluations of semantic embeddings do not correlate with downstream task performance (Chiu et al., 2016; Bakarov, 2018). Similarly, we propose that researchers consider the purpose of personas. If the end goal is replacements for people, even setting aside the significant ethical concerns, they would require very high fidelity. But if personas are used as tools to augment human decision-making in specific contexts, the required fidelity (and even how to measure fidelity) likely varies by task.
@@
There is still (relatively) little research on how best to steer large language models to incorporate new information from prior Agents optimally (Zhang et al., 2024). For example, it is possible that a prior Agent’s response degrades the performance of a future Agent. These questions are highly relevant as practitioners are increasingly using retrieval-augmented generation (RAG) (Gao et al., 2023b). Our package can serve as a useful testbed for researchers who are studying how best to combine and filter new information to complete tasks. In human diffusion, initial behavior has a large effect on cascades (Salganik and Watts, 2008; Muchnik et al., 2013). Plurals can be used to understand: What structures and combination instructions minimize undesirable Agent-based cascades (Ju et al., 2024)?
@@
We chose ANES as an initial persona-generation dataset due to its use in prior work (Argyle et al., 2023b), inclusion of political variables, and updating frequency. Nonetheless, ANES is just one possible generator of data-driven personas and is limited to the United States, does not represent non-citizens, and is heavily focused on demographic and political variables. In future iterations, we plan on adding orthogonal datasets.
@@
Our efficacy studies showed our system is an improvement over zero-shot but this does not necessarily mean it is helpful in general—just that it beats a baseline. We also did not systematically explore the efficacy of Plurals. While vanilla zero-shot and chain-of-thought zero-shot are reasonable baselines since they do not require examples, future work can explore different baselines such as expert-crafted messages, fine-tuned models, or few-shot learning. Second, future work can explore different Plurals configurations. Case Study 3 tested a conventional “focus group” setup, where a Moderator extracts ideas from structured group discussions. Case Studies 4-5 more deeply leveraged Agent interactions, inspired by a mix of (A) “critique-and-revise” Constitutional AI approaches (Bai et al., 2022) and (B) crowdsourced human ideation (Simaei et al., 2023). We simulated a pseudo-crowd to critique-and-revise. We encourage other configurations. Our mechanistic fidelity corresponded to personas and (a slice of) combination instructions. Future work can explore the fidelity of more components. Our steerable moderation case study is a simplified proof of concept since many tasks do not cleanly violate just one principle and not others. Future work can more thoroughly evaluate whether Plurals accurately abstain based on user-defined values. These case studies are a preliminary exploration of Plurals.
Our case studies were limited to political domains. This choice was due to (1) the importance of politics in society, (2) the natural connection between political issues and deliberative democracy, and (3) the feasibility of recruiting group members for validation. However, Plurals can structure interactions among Agents varying in other theoretically-grounded attributes, such as Schwartz’s Theory of Basic Values (Schwartz, 2012), Moral Foundations Theory (Graham et al., 2013), and user types (Bartle, 1996). In future work, our library can be used for other domains (e.g.: education, science, business).